OpenStudy (anonymous):

find dy/dx for the following function y = 9cosx /(1+sinx) u=9cosx v= 1+six u' = -9sinx v' = cosx u'v - v'u / (v)^2 -9sinx (1+sinx)- cosx(9cosx) / (1+sinx)^2 ? not sure if i'm doing this right... answer is -(9/1+sinx) but I'm not getting that.

OpenStudy (anonymous):

You forgot your trig identifty for \[-9\sin ^{2}x-9\cos^{2}x=-9(\sin ^{2}x +\cos ^{2}x)=-9\] Then your top simplifies to: \[-9sinx-9=-9(sinx+1)\] So, one of the sinx+1's in the denominator will cancel with the top one, leaving your answer as required.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You are welcome.